# OCR Hinders RAG: Evaluating the Cascading Impact of OCR on Retrieval-Augmented Generation

### v1

@@ 3. Table Processing:
 - Convert tables to LaTeX.
@@ 5. Output Format:
 Follow these guidelines strictly to ensure accurate and consistent conversion. Only output the OCR results without any extra explanations or comments.
Table S7: Prompt template for VLMs OCR
@@
 You are able to mimic the various errors introduced in OCR recognition. You are given examples with ground truth and OCR results with errors. Your job is to refer to the examples and generate three samples with increasing perturbations based on the user input.
 The new perturbed document could include the following OCR errors:
 1. formatting issues, character recognition errors (like misrecognized letters or numbers), and slight variations in wording.
 2. Redundant content and typos for plain text and formula.
 3. Randomly add and delete table structure controlling character "&" for tabular data.
 4. Randomly add extra columns and rows for tabular data.
 5. Randomly make misaligned columns and rows for tabular data.
 - In the first perturbed document, make mild perturbation with 10% changes.
 - In the second perturbed document, make moderate perturbation with 30% changes.
 - In the third perturbed document, make severe perturbation with 50% changes.
 Output the three perturbed documents within <response></response>.
 You should refer to the following perturbed example to generate the perturbed document.
 Example ground truth: {gt_sample}
 Example results with mild perturbation: {mild_sample}
 Example results with moderate perturbation: {moderate_sample}
 Example results with severe perturbation: {severe_sample}
User:
 <Begin Document>{gt_structured_data}<End Document>
Table S8: Prompt template for Semantic Noise introduction
Retrieval Generation Overall
[…]
Simple Prompt
Qwen2-VL-72B 52.50 72.27 26.29 40.65 23.09 35.30
InternVL2-Llama3-76B 41.23 57.31 20.07 32.36 20.03 30.40
Detailed Prompt
Qwen2-VL-72B 53.16 72.97 26.72 41.23 23.45 35.91
InternVL2-Llama3-76B 42.43 57.51 20.74 32.89 20.58 31.23
Table S9: Effects of Different OCR Prompts on VLMs for OCR. The detailed prompt is used due to its consistently superior performance.
Law Finance Textbook Manual News Academia Ground Truth LCS@1 64.31 37.96 79.89 74.18 72.89 56.92 LCS@5 90.08 70.97 86.14 94.28 87.34 85.59 MinerU LCS@1 59.01 31.22 55.75 58.92 58.31 48.27 LCS@5 83.50 58.77 62.06 80.90 76.15 70.61 Marker LCS@1 61.34 36.91 60.55 67.31 65.80 49.76 LCS@5 86.20 67.21 65.03 89.23 83.15 74.08 GOT LCS@1 52.18 31.70 49.28 62.00 4.08 48.69 LCS@5 74.69 59.49 66.52 84.09 6.74 73.20 Nougat LCS@1 56.36 24.11 24.11 47.67 0.60 49.01 LCS@5 79.30 44.96 44.96 62.29 1.03 70.91 Qwen2-VL-72B LCS@1 58.21 34.52 62.68 64.86 54.17 39.23 LCS@5 82.36 62.84 66.60 87.24 67.25 57.06 InternVL2 -Llama3-76B LCS@1 41.31 17.20 61.26 66.44 17.33 42.02 LCS@5 59.50 32.07 65.51 87.83 23.54 59.86
Table S10: Comparison of OCR solutions on various domains in the retrieval stage using LCS@1 and LCS@5 metrics. Bold indicates the best performance, and underline indicates the second-best performance.
Law Finance Textbook Manual News Academia Ground Truth F1 45.85 35.25 45.85 50.24 40.38 41.60 EM 30.25 25.63 30.25 30.42 23.28 30.24 MinerU F1 43.95 30.53 24.34 44.14 37.02 41.09 EM 28.85 21.89 11.67 26.13 22.02 30.98 Marker F1 43.50 31.50 21.56 47.16 36.41 36.27 EM 27.95 21.89 9.86 27.73 21.98 26.13 GOT F1 39.69 31.06 24.69 44.70 5.55 37.28 EM 25.82 22.15 11.85 26.15 3.23 27.04 Nougat F1 41.36 26.51 20.02 37.38 2.68 37.25 EM 26.73 18.82 9.00 21.35 1.11 27.65 Qwen2-VL-72B F1 43.13 34.69 23.08 46.80 29.63 34.04 EM 27.57 24.46 10.73 27.97 16.13 24.58 InternVL2 -Llama3-76B F1 34.89 25.62 21.05 44.85 12.18 33.17 EM 22.76 18.36 10.15 26.26 6.14 24.15
Table S11: Comparison of OCR solutions on various domains in the generation stage using F1-score and EM metrics. Bold indicates the best performance, and underline indicates the second-best performance.
Law Finance Textbook Manual News Academia Ground Truth F1@1 44.35 33.10 28.90 47.02 39.97 40.59 EM@1 30.84 24.49 14.99 28.55 23.20 30.09 F1@5 47.36 37.40 25.09 53.46 40.79 42.61 EM@5 29.65 26.77 10.47 32.29 23.35 30.39 MinerU F1@1 41.44 29.16 14.99 40.71 37.07 41.52 EM@1 28.43 20.86 14.13 24.22 22.48 32.25 F1@5 46.45 31.90 21.90 47.58 36.96 40.66 EM@5 29.27 22.93 9.21 28.03 21.55 29.71 Marker F1@1 41.97 29.89 23.79 45.26 35.90 36.02 EM@1 23.63 21.26 11.89 26.87 22.99 26.64 F1@5 45.02 33.09 19.32 49.05 36.92 36.53 EM@5 27.52 22.52 7.84 28.59 20.98 25.64 GOT F1@1 38.47 29.37 24.67 41.48 5.47 37.05 EM@1 26.32 21.10 12.14 24.37 3.30 27.69 F1@5 40.92 32.74 24.71 47.92 5.64 37.51 EM@5 25.31 23.20 11.56 27.92 3.16 26.39 Nougat F1@1 39.95 24.57 21.59 35.13 2.61 37.74 EM@1 27.29 17.60 10.15 20.47 1.15 28.84 F1@5 42.78 28.45 18.46 39.64 2.76 36.76 EM@5 26.17 20.04 7.84 22.23 1.07 26.47 Qwen2-VL-72B F1@1 40.80 32.05 32.05 44.30 29.62 34.61 EM@1 27.31 22.98 22.98 26.98 17.17 26.11 F1@5 45.45 37.33 20.73 49.29 29.64 33.47 EM@5 27.82 25.95 8.63 28.96 15.09 23.06 InternVL2 -Llama3-76B F1@1 23.63 24.53 24.86 44.04 12.14 35.15 EM@1 23.63 17.66 12.58 26.92 6.54 26.39 F1@5 35.35 26.71 17.23 45.66 12.22 31.19 EM@5 21.88 19.06 7.73 25.60 5.75 21.91
Table S12: Comparison of OCR solutions on various domains in overall performance. The performance is presented with top-1 (EM@1, F1@1) and top-5 (EM@5, F1@5) retrieved chunks. Bold indicates the best performance, and underline indicates the second-best performance.
Figure S1: The impact of Semantic Noise ([S] dashed line) and Formatting Noise ([F] solid line) on different retrievers. The horizontal axis represents the edit distance, and higher values indicate greater OCR noise. We report the LCS@1 of each retriever on questions from all evidence sources.
Figure S2: The impact of Semantic Noise ([S] dashed line) and Formatting Noise ([F] solid line) on different LLMs. The horizontal axis represents the edit distance, and higher values indicate greater OCR noise. We report the F1-score of each LLM on questions with all evidence sources.
Figure S3: The impact of Semantic Noise ([S] dashed line) and Formatting Noise ([F] solid line) on different RAG systems, using top-1 in the retrieval stage. The horizontal axis represents the edit distance, and higher values indicate greater OCR noise. Overall performance is presented for top-1 (F1@1) retrieved chunks.
Figure S4: One of the real table cases used to guide the introduction of Semantic Noise. The upper left is the original table in the ground truth, and the upper right is a real example from the OCR result of MinerU. The lower left and lower right are the results of moderate and severe perturbation to the original table after using the real example as guidance. For a better show, we modified some latex codes manually so that most part of the table structure can be displayed normally.

### v2

@@ 3. Table Processing:
 - Convert all tables to LaTeX format.
[…]
 - Convert all Charts to LaTeX format.
@@ 5. Output Format:
 Please strictly follow these guidelines to ensure accuracy and consistency in the conversion. Your task is to accurately convert the content of the PDF image using these format requirements without adding any extra explanations or comments.
Table S11: Complex prompt for VLMs OCR
@@
 You are an AI specialized in document question-answering verification. Your mission is to analyze the given question-answering pairs and follow the instructions. Your response must be true and accurate, and no additional content should be output.
1. Question type check
 Dose the question match the task description: {detailed_task_description}
 Make sure the question meets the required task context.
2. Evidence relevance Check
 Dose the provided evidence context relate to the question provided? Does the answer accurately reflect the information in the evidence context? Ensure the question is formulated based on information explicitly stated. The question should not introduce concepts unrelated to the document’s content.
3. Clarity and Precision
 Is the question clear and unambiguous? And is the answer concise and precise? Ensure the language is straightforward and easily understandable, and avoid complex phrasing that may confuse the reader.  The intention of the question and answer pair must be clear and direct, avoiding verbosity and unnecessary detail.  Ensure the answer fully addresses the question without omitting crucial information.
{qas}
Table S12: Q&A Verification Prompt
V Case Study
Fig. S2 to  Fig. S11 show some cases of GOT, MinerU, and Qwen2.5VL-72B on OHRbench. For each case, we indicate the evidence source and answer, giving the OCR result of different models and the responses at the retrieval and generation stages.
Figure S2: A case using text as the evidence source on a distorted academic document.
Figure S3: A case using formula as the evidence source on a scanned academic document.
Figure S4: A case using text in multi-pages as the evidence source on an academic document.
Figure S5: A case using formula as the evidence source on a scanned textbook.
Figure S6: A case using text as the evidence source on a handwritten textbook.
Figure S7: A case using text as the evidence source on a handwritten textbook.
Figure S8: A case using table as the evidence source on a financial report.
Figure S9: A case using table as the evidence source on a scanned academic paper.
Figure S10: A case using text with reading order as the evidence source on a scanned newspaper.
Figure S11: A case using text as the evidence source on a distortion manual.